Please help solve by substitution method.

y = x + 4
3y -5x = 6

Answer by erikaii.19(2)

y = x + 4
3y -5x = 6

a) 3y-5x=6
3(x+4)-5x=6
3x+12-5x=6
-2x=6-12
-2x=-6
x=3

b) y=x+4
y=3+4
y=7



CHECKING

1) 7=3+4
7=7

2) 3(7)-5(3)=6
21-15=6
6=6

:)